OpenStudy (anonymous):

2y – 6x = -2 so y = 3x-1 y intercept is (0,-1) and slope is 3 But how do I find the second point on the line using slope please ?

OpenStudy (luigi0210):

Using slope? Well\[\Large slope = \frac{rise}{run}=\frac{3}{1}\]

OpenStudy (luigi0210):

Rise= y Run =x

OpenStudy (anonymous):

thanks but i still don't know what they did to get the second point of (3,-2) if you can let me know, im sure im just blind :)

OpenStudy (luigi0210):

Hmm, you sure it's (3, -2)?

OpenStudy (anonymous):

oh.... no, i've gone wrong there !

OpenStudy (anonymous):

(1,2)

OpenStudy (luigi0210):

There that makes more sense :) Okay, so you know how the x intercept is \[\large (0, -1) => (x, y)\] right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

that was y intercept ?

OpenStudy (anonymous):

sorry, x intercept :)

OpenStudy (luigi0210):

Based on this point, what they did was go up 3 and over 1. \[\Large slope = \frac{rise}{run}=\frac{3}{1}=\frac{y}{x}\]  So \[\Large (0(x), -1(y))+slope~3(\frac{3}{1}(\frac{y}{x}))=([0+1], [-1+3])=(1, 2)\]